OpenStudy (anonymous):

Can I get some help fan and medal.... Rationalize the denominator of square root of negative four over open parentheses 7 minus 3 I close parentheses plus open parentheses 2 plus 5 i. quantity of negative 16 plus 18 I over 7 quantity of 16 plus 18 I over 145 quantity of 4 plus 18 I over 85 quantity of negative 4 plus 18 I over 77

OpenStudy (tkhunny):

The best way to get help is by showing your won work, not by spamming with multiple tags. Give it a go! Let's see what you get.

OpenStudy (anonymous):

I have literally no idea lol, I know it's something to do with conjugates though

OpenStudy (anonymous):

:/

OpenStudy (tkhunny):

You simply can't have "no idea". Did you sleep through class? Do you have a book? Are you in the wrong class? To "rationalize" is to find the right magic number that will get rid of everything that isn't rational. \(\sqrt{2}\) Well, you need another \(\sqrt{2}\) \(1 + \sqrt{3}\) Well, you need the conjugate, \(1 - \sqrt{3}\), takign advantage of the idea that (a+b)(a-b) = a^2 - b^2 \(3 + 2i\) Well, you need the complex conjugate \(3 - 2i\), for the same reason as the previous example. Rationalize the denominator of square root of negative four over open parentheses 7 minus 3 I close parentheses plus open parentheses 2 plus 5 i. Is it this? \(\dfrac{\sqrt{-4}}{(7-3i)+(2+5i)}\)?

OpenStudy (anonymous):

That is the equation yes, now can u walk me through step by step I read the lesson Online school, but it wasn't very detailed only 1 page long

OpenStudy (tkhunny):

Just add the stuff in the denominator. Add the like terms and simplify it.

OpenStudy (anonymous):

So the denominator becomes 9+2i?

OpenStudy (tkhunny):

Yes. Now, supply the Complex Conjugate of that number.

OpenStudy (anonymous):

9-2i?

OpenStudy (tkhunny):

Perfect. Multiply both numerator and denominator by that value.

OpenStudy (anonymous):

thats where i have trouble @tkhunny

OpenStudy (tkhunny):

So? Do it!!! Did you understand this part: Taking advantage of the idea that (a+b)(a-b) = a^2 - b^2 That's all we're doing. ( 9+2i)( 9 -2i) = ??? Multiply away, just like any other binomial.

OpenStudy (anonymous):

-4+18i? @tkhunny

OpenStudy (tkhunny):

What? I know you have multiplied binomials, before. ( 9+2i)( 9 -2i) = 91 + 18i - 18i - 4i^2 Ring any bells?

OpenStudy (anonymous):

yeah I got that, then what about simplifying it?

OpenStudy (tkhunny):

You tell me. Add the like terms and watch the magic.

OpenStudy (anonymous):

yeah I get 91-4i2 (squared) but thats not in the choices for denominator or numerator

OpenStudy (tkhunny):

Whoops! That's 81. Sorry for the typo. Okay, now what was the point of the Complex Conjugate? Were we just making rice pudding or did we have a purpose?

OpenStudy (anonymous):

Seems like rice pudding still :/

OpenStudy (tkhunny):

It may SEEM like that. But, what was our purpose? We wanted to accomplish something. What was it?

OpenStudy (anonymous):

Weren't we gonna multiply by numerator and denominator?

OpenStudy (tkhunny):

Well, okay, that was the method, but that was not the purpose. We wanted to clear up the denominator and get all the Imaginary stuff out of there. Do you recall this purpose?

OpenStudy (anonymous):

Sorry I am just so lost on this part of the lesson, I recall the purpose now.

OpenStudy (tkhunny):

We have 81 - 4i^2 Does that look Real and Rational?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

?? What's that "i" doing in there? That's not Real, is it?

OpenStudy (anonymous):

oh wait woops that's imaginary so how do we get rid of it?

OpenStudy (tkhunny):

i^2 = What???

OpenStudy (anonymous):

-i? or -1?

OpenStudy (tkhunny):

\(\sqrt{-1} = i\) therefore, \(i^{2} = -1\) 81 - 4i^2 = 81 - 4(-1) = 81 + 4 = 85 Wow!! Anyone remember where this all started? \(\dfrac{\sqrt{-4}}{(7-3i)+(2+5i)} = \dfrac{i\sqrt{4}}{(9+2i)} = \dfrac{2i}{(9+2i)}\) That was just simplifying. Then, we introduced the Complex Conjugate. \(\dfrac{2i}{(9+2i)}\cdot\dfrac{9-2i}{9-2i} = \dfrac{18i - 4i^{2}}{81-4i^{2}} = \dfrac{18i + 4}{81+4} = \dfrac{4+18i}{85}\) It wasn't nearly as bad all in one place without all the discussion and missteps. Follow each step and make sure you know where it came from.

OpenStudy (anonymous):

okk thanks, I was lost because the first step after "then we introduced the complex conjugate" I didn't have and such a mess, but I totally understand, thanks so much @tkhunny

OpenStudy (tkhunny):

You really hung in there. Good work! Now, go do 100 more! :-)

OpenStudy (anonymous):

lol, I'll try.